OpenStudy (anonymous):

Tom’s uncle owns a triangular piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one-half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side? I came up with this equation but I don't know if it is right or not.:P = 2L + 2W

OpenStudy (lgbasallote):

let x = longest side shortest side = 30 + x/2 second longest side = x - 2 perimeter is sum of these so... x + 30 + x/2 + x - 2 = 378 2x + 28 + x/2 = 378 5x/2 + 28 = 378 5x/2 = 350 5x = 700 x = 140 can you do the rest?

OpenStudy (lgbasallote):

tell me if there's something you don't get there

OpenStudy (anonymous):

did my equation that I came up with show on here? I don't see it. I wanted to knwo if I was on the right track

OpenStudy (lgbasallote):

what you wrote was perimeter of a quadrilateral or commonly..rectangles this is perimeter of triangle

OpenStudy (anonymous):

378=2X+2+ 2A=30/2X

OpenStudy (lgbasallote):

the equation i wrote is the equation you need...i see no other equation applicable

OpenStudy (anonymous):

ok